### v2

# The dual of Philo’s shortest line segment problem

Abstract.
We study the dual of Philo’s shortest line segment problem which asks to find the optimal line segments passing through two given points, with a common endpoint, and with the other endpoints on a given line. The provided solution uses multivariable calculus and geometry methods. Interesting connections with the angle bisector of the triangle are explored. A similar property of a symedian of a triangle is conjectured.
Key words and phrases:
Philo’s line, geometric inequalities, geometric maxima and minima, angle bisector, symedian.
1991 Mathematics Subject Classification:
Primary 51M16, 51M25; Secondary 51M04, 52A38, 52A40.
1. Introduction
Maximum-minimum problems in geometry played crucial role in the history of development of mathematical methods [11], Chapter VII, [8], [21] Chapters VIII-XIX. Many attempts to solve variety of geometry problems for optimization lead to the develoment of differential and integral calculus and later to the calculus of variations. Maxima-minima problems also play important role in the pedagogy of mathematics [3], [25]. In mathematical olympiads for schools or universities, and journals with problem solving columns such as Crux Mathematicorum and American Mathematical Monthly one can often find at least one geometric inequality included in its list of proposed problems. Recent breakthough in automation of proofs for classical euclidean geometry theorems is yet to be done for the geometric inequalities [27]. Mathematicians noticed very early that some of the geometric optimization problems do not have solutions and even when they do have a solution it is not always possible to construct it with euclidean instruments (a compass and a straightedge). The problem about minimal line segment with fixed point and inscribed into a given angle is one of them [9], p. 405-406, [23], p. 36-37, [24] p. 33, 179-181, [18], [5], [19]. In this problem an angle $ABC$ and a point $D$ in it is given and it is asked to find the line $EF$ through this point which has shortest line segment $EF$ cut by this angle. Sir I. Newton mentioned this problem as an example for his method of fluxions (see [20], p. 45, Problem 5.III). The solution of this problem is known in the literature as Philo’s line [13], [14] p. 39 and 234-238, [29] p. 182-183, [28], [16] p. 115–116. If $BG$ is the perpendicular to $EF$ from point $B$, then $EF$ is shortest when $|ED|=|FG|$ (see Figure 1).
Figure 1. Philo’s line and the dual problem
Philo or Philon of Byzantium found its connection with the classical problem of the dublication of cube [7], p. 107, [31], [22] p. 24, [26]. The digits of the length of Philo’s line through the incenter of triangle with sidelengths 3,4,5 appear in the OEIS as A195284 [17]. Generalizations for non-euclidean geometry and higher dimensions appeared in [12], [30]. Related problem of the longest or shortest chords of conics and curves in general have similar solutions [4], [6], [2].
In projective geometry the principle of duality is used to explain the similarity between theorems where the roles of lines and points are interchanged. If we use this idea in the above extremum problem, then we obtain the following dual problem ([2], p. 4). Given line $l$ and points $A$ and $B$ find point $C$ with ray $CA$ and $CB$ intersecting line $l$ at $D$ and $E$, respectively, such that the sum of lengths of $CD$ and $CE$ is minimal (see Figure 1 and Table 1). This problem is surprisingly much more challenging than Philo’s shortest line problem. In the following text this problem is solved using combination of methods from calculus and geometry. At the end of the paper we mentioned other interesting formulae about the configuration and several open problems.
Philo’s line problem Dual problem
Fixed point(s) 1 ($D$) 2 ($A$ and $B$)
Fixed line(s) 2 ($BA$ and $BC$) 1 ($l$)
Moving point(s) 2 ($E$ and $F$) 1 ($C$)
Moving line(s) 1 ($EF$) 2 ($CD$ and $CE$)
Table 1. The number of elements in Philo’s line problem and its dual.
2. Main results.
As was mentioned before the optimal solution of geometric extrema problems can sometimes be non-constructible with the help of a compass, a ruler, and the given data. Without the proof of the existence of the optimal solution the proof is incomplete. For example, in Hadamard’s classical geometry problem book some of such extrema problems are provided with a comment like "assume that a maximum exists" (see e.g. Exercise 366 in p. 313, Exercise 418b in p. 322 of [15]). Most of the time it is possible to refer to calculus results for more rigorous proof but it makes the solution less appealing for geometrically inclined mind. Another possible way out from this situation is to give or construct the optimal solution first in the diagram and then ask to prove that the other possibilities are not optimal. In the problem about Philo’s line we can give $\triangle BEF$ and points $D$ and $G$ on its side $EF$ such that $|ED|=|FG|$, $BG\perp EF$, and then prove that any other line segment $E_{1}F_{1}$ through point $D$ has a greater length (see Figure 1). We can formulate the mentioned dual problem above in the same way by giving the optimal solution first. It will be clear from the proof why it is optimal and how it was found in the first place.
Theorem 2.1.
Let $CDE$ be a triangle with acute angles at vertices $D$ and $E$. Let $CK$ be the angle bisector of triangle $CDE$. Drop perpendiculars $KL$ and $KM$ to sides $CD$ and $CE$, respectively. Take points $A$ and $B$ on sides $CD$ and $CE$, respectively, such that $|AC|=|LD|$, $|BC|=|ME|$. Then for any point $C_{1}$ different from $C$, such that rays $C_{1}A$ and $C_{1}B$ intersect line $DE$ at points $D_{1}$ and $E_{1}$, respectively, inequality $|C_{1}D_{1}|+|C_{1}E_{1}|>|CD|+|CE|$ holds true (see Figure 2).
Figure 2. Inequality $|C_{1}D_{1}|+|C_{1}E_{1}|>|CD|+|CE|$
Proof.
Let us denote the coordinates of points as $A(a,b),\ B(c,d),\ D_{1}(x_{1},0)$ and $E_{1}(x_{2},0)$. Without loss of generality we can assume that $a<c$ and $b\geq d>0$. If we denote coordinates of $C_{1}$as $(x,y)$, then $x_{1}=\frac{ay-bx}{y-b}$ and $x_{2}=\frac{cy-dx}{y-d}$. Here $(x,y)\in\mathbb{D}$, where $\mathbb{D}=\{(x,y)|x\in(-\infty,+\infty),\ y>b\}$. The last inequality is needed to guarantee the condition in the satatement of the problem which says that rays $C_{1}A$ and $C_{1}B$ intersect line $DE$. We obtain that
$\left|C_{1}D_{1}\right|+\ \left|C_{1}E_{1}\right|=\sqrt{\left(x-x_{1}\right)^{2}+y^{2}}+\sqrt{\left(x-x_{2}\right)^{2}+y^{2}}$
$=y\left(\sqrt{\left(\frac{x-a}{y-b}\right)^{2}+1}+\sqrt{\left(\frac{x-c}{y-d}\right)^{2}+1}\right)=:f(x,y).$
Note that $f(x,y)$ is continuous in $\mathbb{D}$ and $f(x,y)\rightarrow+\infty$ as $x\rightarrow\pm\infty$, $y\rightarrow b$ when $x\neq a$. Similarly, $f(x,y)\rightarrow+\infty$ as $y\rightarrow+\infty$. Also $\lim_{(x,y)\rightarrow(a,b)}f(x,y)$ does not exist and depends on the direction of approach to point $(a,b)$. Indeed, if $b>d$, then
$\lim_{t\rightarrow 0^{+}}f(a+k_{1}t,b+k_{2}t)=b\left(\sqrt{\frac{k_{1}^{2}+k_{2}^{2}}{k_{2}^{2}}}+\frac{\sqrt{(b-d)^{2}+\left(a-c\right)^{2}}}{b-d}\right),$
where $k_{2}>0$. This limit is minimal when $k_{1}=0$ which corresponds to the direction of line $x=a$. If $b=d$, then the last limit is also infinite. Function $f(x,y)$ is also bounded from below because it is a positive function. Then either there is a point $(x^{*},y^{*})\in\mathbb{D}$ where function $f(x,y)$ is minimal or this function approaches its infimum when $(x,y)\rightarrow(a,b)$ $(y>b)$. We will study the last case later. In the first case $f^{\prime}_{x}(x^{*},y^{*})=0$, $f^{\prime}_{y}(x^{*},y^{*})=0$. Obviously, $a\leq x^{*}\leq c$, because if $x<a$, then $f(x,y)>f(a,y)$, and if $x>c$, then $f(x,y)>f(c,y)$. Note also that if $y$ is fixed, then function $f(x,y)$ is convex with respect to $x$. Indeed,
$f_{xx}=\frac{y(y-b)}{\sqrt{{\left(\left(x-a\right)^{2}+(y-b)^{2}\right)^{3}}}}+\frac{y\left(y-d\right)}{\sqrt{{\left(\left(x-c\right)^{2}+(y-d)^{2}\right)^{3}}}}>0.$
We calculate
$f_{x}(a,y)=\frac{y\left(a-c\right)}{\left(y-d\right)\sqrt{{\left(y-d\right)^{2}+\left(a-c\right)^{2}}}}<0,$
$f_{x}(c,y)=\frac{y\left(c-a\right)}{\left(y-b\right)\sqrt{{\left(y-b\right)^{2}+\left(a-c\right)^{2}}}}>0.$
Therefore $a<x^{*}<c$. In order to locate point $x^{*}$, note that equality $f^{\prime}_{x}(x,y)=0$ implies
$\frac{x-a}{(y-b)^{2}\sqrt{\left(\frac{x-a}{y-b}\right)^{2}+1}}=\frac{c-x}{(y-d)^{2}\sqrt{\left(\frac{x-c}{y-d}\right)^{2}+1}}\Leftrightarrow$
$\frac{x-a}{(y-b)\sqrt{(x-a)^{2}+(y-b)^{2}}}=\frac{c-x}{(y-d)\sqrt{(x-c)^{2}+(y-d)^{2}}}.$
Figure 3. $|K_{1}L_{1}|=|K_{2}M_{1}|.$
The last equality can be interpreted geometrically in the following way. Take points $L_{1}$ and $M_{1}$ on sides $C_{1}D_{1}$ and $C_{1}E_{1}$, respectively, such that $|AC_{1}|=|L_{1}D_{1}|$, $|BC_{1}|=|M_{1}E_{1}|$. Take points $K_{1}$ and $K_{2}$ on line $D_{1}E_{1}$ such that $K_{1}L_{1}\perp C_{1}D_{1}$ and $K_{2}M_{1}\perp C_{1}E_{1}$. Equality $f^{\prime}_{x}(x,y)=0$ is equivalent to say that $|K_{1}L_{1}|=|K_{2}M_{1}|$ (see Figure 3). Indeed, if we draw the line through $C_{1}$ and parallel to $D_{1}E_{1}$ and take points $F$ and $G$ such that $AF\perp AC_{1}$ and $BG\perp BC_{1}$, then $|AF|=|K_{1}L_{1}|$ and $|BG|=|K_{2}M_{1}|$. On the other hand,
$|AF|=|AC_{1}|\cdot\tan{\angle AC_{1}F}=\sqrt{(x-a)^{2}+(y-b)^{2}}\cdot\frac{y-b}{a-x}.$
Similarly,
$|BG|=\sqrt{(x-c)^{2}+(y-d)^{2}}\cdot\frac{y-d}{c-x}.$
Therefore
$|K_{1}L_{1}|=|K_{2}M_{1}|.$ (1)1( 1 )
The other condition ($f^{\prime}_{y}(x,y)=0$) can also be interpreted geometrically but its little harder than the previous condition. First, drop perpendicular $C_{1}H$ to side $D_{1}E_{1}$. Next, we calculate
$f^{\prime}_{y}(x,y)=\left(\sqrt{\left(\frac{x-a}{y-b}\right)^{2}+1}+\sqrt{\left(\frac{x-c}{y-d}\right)^{2}+1}\right)+$
$y\left(-\frac{\left(x-a\right)^{2}}{\sqrt{1+\frac{\left(x-a\right)^{2}}{\left(y-b\right)^{2}}}\,\left(y-b\right)^{3}}-\frac{\left(x-c\right)^{2}}{\sqrt{1+\frac{\left(x-c\right)^{2}}{\left(y-d\right)^{2}}}\,\left(y-d\right)^{3}}\right).$
If $a<x<c$, then equality $f^{\prime}_{y}(x,y)=0$ can be interpreted as
$|C_{1}D_{1}|+|C_{1}E_{1}|=\frac{|D_{1}H|^{2}}{|L_{1}D_{1}|}+\frac{|E_{1}H|^{2}}{|M_{1}E_{1}|}.$ (2)2( 2 )
Indeed, first note that
$\frac{1}{y}\left(|C_{1}D_{1}|+|C_{1}E_{1}|\right)=\sqrt{\left(\frac{x-a}{y-b}\right)^{2}+1}+\sqrt{\left(\frac{x-c}{y-d}\right)^{2}+1}.$
On the other hand note that
$\frac{\left(x-a\right)^{2}}{\sqrt{1+\frac{\left(x-a\right)^{2}}{\left(y-b\right)^{2}}}\,\left(y-b\right)^{3}}+\frac{\left(x-c\right)^{2}}{\sqrt{1+\frac{\left(x-c\right)^{2}}{\left(y-d\right)^{2}}}\,\left(y-d\right)^{3}}$
$=\frac{\frac{\left(x-a\right)^{2}}{\left(y-b\right)^{2}}}{\sqrt{{\left(x-a\right)^{2}}+{\left(y-b\right)^{2}}}}+\frac{\frac{\left(x-c\right)^{2}}{\left(y-d\right)^{2}}}{\sqrt{{\left(x-c\right)^{2}}+{\left(y-d\right)^{2}}}}$
$=\frac{\tan^{2}{\angle AC_{1}F}}{|AC_{1}|}+\frac{\tan^{2}{\angle BC_{1}G}}{|BC_{1}|}.$
We know that $\angle AC_{1}F=\angle D_{1}$, $\angle BC_{1}G=\angle E_{1}$, $|AC_{1}|=|L_{1}D_{1}|$ and $|BC_{1}|=|M_{1}E_{1}|$. Therefore,
$|C_{1}D_{1}|+|C_{1}E_{1}|=|C_{1}H|^{2}\left(\frac{\tan^{2}{\angle D_{1}}}{|L_{1}D_{1}|}+\frac{\tan^{2}{\angle E_{1}}}{|M_{1}E_{1}|}\right),$
Figure 4. One possible order of points on line $D_{1}E_{1}$.
from which it follows that (2) holds true. Since $\frac{|D_{1}H|^{2}}{|L_{1}D_{1}|}=|D_{1}H|\cdot\frac{|C_{1}D_{1}|}{|D_{1}K_{1}|}$ and $\frac{|E_{1}H|^{2}}{|M_{1}E_{1}|}=|E_{1}H|\cdot\frac{|C_{1}E_{1}|}{|E_{1}K_{2}|}$, equality (2) can also be written as
$|C_{1}D_{1}|+|C_{1}E_{1}|=|C_{1}D_{1}|\cdot\frac{|D_{1}H|}{|D_{1}K_{1}|}+|C_{1}E_{1}|\cdot\frac{|E_{1}H|}{|E_{1}K_{2}|},$
which we can rewrite as
$|C_{1}D_{1}|\cdot\left(1-\frac{|D_{1}H|}{|D_{1}K_{1}|}\right)=|C_{1}E_{1}|\cdot\left(\frac{|E_{1}H|}{|E_{1}K_{2}|}-1\right),$ (3)3( 3 )
or finally as
$|C_{1}D_{1}|\cdot\frac{|K_{1}H|}{|D_{1}K_{1}|}=|C_{1}E_{1}|\cdot\frac{|K_{2}H|}{|E_{1}K_{2}|}.$
From this we obtain
$\frac{|D_{1}C_{1}|}{|C_{1}E_{1}|}\cdot\frac{|E_{1}K_{2}|}{|K_{2}H|}\cdot\frac{|HK_{1}|}{|K_{1}D_{1}|}=1.$ (4)4( 4 )
If $AU$ is the angle bisector of triangle $D_{1}C_{1}E_{1}$, then by the angle bisector theorem $\frac{|D_{1}C_{1}|}{|C_{1}E_{1}|}=\frac{|D_{1}U|}{|UE_{1}|}$. By noting this we can write (4) as
$\frac{|D_{1}U|}{|UE_{1}|}\cdot\frac{|E_{1}K_{2}|}{|K_{2}H|}\cdot\frac{|HK_{1}|}{|K_{1}D_{1}|}=1.$ (5)5( 5 )
If we denote $|D_{1}H|=v$, $|K_{2}H|=w$, $|K_{2}U|=z$, $|K_{1}U|=t$, $|K_{1}E_{1}|=u$ (see Figure 4), equation (5) can be written simply as
$\frac{v+w+z}{t+u}\cdot\frac{u+t+z}{w}\cdot\frac{w+z+t}{v+w+z+t}=1,$
which after simplifications gives
$t^{2}v+t^{2}z+tuv+tuz+2tvz+2twz+2tz^{2}$
$+uvz+uwz+uz^{2}+vwz+vz^{2}+w^{2}z+2wz^{2}+z^{3}$
$=t(tv+tz+uv+uz+2vz+2wz+2z^{2})$
$+z(uv+uw+uz+vw+vz+w^{2}+2wz+z^{2})=0.$
It is possible only when $z=t=0$, which means that points $K_{1},K_{2},$ and $U$ coincide. The other cases when points $K_{1}$ or $K_{2}$ are outside of the side $D_{1}E_{1}$, or when point $K_{1}$ is closer to point $D_{1}$ than point $K_{2}$ can be considered similarly. For example, if the order of points is as in Figure 5, where the notations for the lengths of the line segments has slightly changed, then
$\frac{v+w+t}{z+u}\cdot\frac{u}{z+t+w}\cdot\frac{w}{v+w}=1,$
which is not possible. Indeed, in this case since $\frac{v+w+t}{z+u}=\frac{|D_{1}C_{1}|}{|C_{1}E_{1}|}<1,$ $v+w+t<z+u$, and since
$u=|K_{2}M_{1}|\cdot\csc{\angle E_{1}}=|K_{1}L_{1}|\cdot\csc{\angle E_{1}}<|K_{1}L_{1}|\cdot\csc{\angle D_{1}}=v+w,$
$u<v+w$, and therefore
$\frac{v+w+t}{z+u}\cdot\frac{u}{v+w}\cdot\frac{w}{z+t+w}<1.$
Figure 5. Another order of points on line $D_{1}E_{1}$.
Finally, if the order of points is as in Figure 6, where again the notations for the lengths of the line segments has slightly changed, then
$\frac{w+v+z}{t}\cdot\frac{t+z}{v}\cdot\frac{v+z+t+u}{w+v+z+t+u}=1,$
which again simplifies to an impossible equality
$t^{2}w+t^{2}z+tuw+tuz+2tvz+2twz+2tz^{2}$
$+uvz+uwz+uz^{2}+v^{2}z+vwz+2vz^{2}+wz^{2}+z^{3}=0.$
Figure 6. One more order of points on line $D_{1}E_{1}$.
Note that because of equality (1), we only need to consider the cases when point $U$ is between points $K_{1}$ and $K_{2}$. Otherwise, if $U$ is not between points $K_{1}$ and $K_{2}$, then one can drop perpendiculars from point $U$ to sides $C_{1}D_{1}$ and $C_{1}E_{1}$ and one of them will be larger than the other one, which is impossible. Similarly, we do not need to consider the cases when points $K_{1}$ and $K_{2}$ are on different sides of point $H$, because in these cases one of the sides of equality (3) is positive while the other side is negative and therefore (3) does not hold.
If $\angle D_{1}=\frac{\pi}{2}$ ($x=a$), then point $K_{1}$ goes to infinity and equation (5) simplifies to $\frac{|D_{1}U|}{|UE_{1}|}\cdot\frac{|E_{1}K_{2}|}{|K_{2}H|}=1,$ which simply means that points $K_{2}$ and $U$ coincide. Similarly, if $\angle E_{1}=\frac{\pi}{2}$ ($x=c$), then point $K_{2}$ goes to infinity and equation (5) simplifies to $\frac{|D_{1}U|}{|UE_{1}|}\cdot\frac{|HK_{1}|}{|K_{1}D_{1}|}=1,$ which implies that points $K_{1}$ and $U$ coincide.
We proved that if $C_{1}(x^{*},y^{*})$ is a minimum point of $f(x,y)$, then endpoint $U$ of angle bisector $C_{1}U$ coincides with points $K_{1},K_{2}$. Note that triangle $CDE$ and its angle bisector $CK$ satisfy all these conditions together with (1) and (2). In order to complete the proof, we need to show that there are no minimum points other than $C$.
We will now study the locus of intersection point of line $L_{1}K_{1}$ and $M_{1}K_{2}$ when $|AD_{1}|=|BE_{1}|$ and $D_{1}(x_{1},0)$ changes on line $DE$, and then intersection of this locus with line $DE$ (see Figure 7). If $b=d$, then by symmetry this locus is the perpedicular bisector of the line segment $AB$, which always intersect line $DE$ at a unique point. So, we can assume that $b>d$. Let us denote $|AD_{1}|=|BE_{1}|=t$. Then $x_{1}=a-\sqrt{t^{2}-b^{2}}$ and $x_{2}=c+\sqrt{t^{2}-d^{2}}$. Then the equations of lines $AD_{1}$ and $BE_{1}$ are $\frac{y-b}{x-a}=\frac{-b}{x_{1}-a}$ and $\frac{y-d}{x-c}=\frac{-d}{x_{2}-c}$, respectively. The coordinates of intersection point $C_{1}(x_{0},y_{0})$ of these lines are
${x_{0}}=-\frac{d\sqrt{t^{2}-b^{2}}\,\sqrt{t^{2}-d^{2}}+bd^{2}-bt^{2}}{t^{2}\left(b^{2}-d^{2}\right)\left(d^{2}-t^{2}\right)}\times$
×$\times\left(\left(-cd\sqrt{t^{2}-d^{2}}-\left(b-d\right)\left(d^{2}-t^{2}\right)\right)\sqrt{t^{2}-b^{2}}+ab\left(d^{2}-t^{2}\right)\right),$
${y_{0}}=\frac{db\left(\sqrt{t^{2}-d^{2}}\,b-d\sqrt{t^{2}-b^{2}}\right)\left(\sqrt{t^{2}-d^{2}}+\sqrt{t^{2}-b^{2}}-a+c\right)}{t^{2}\left(b^{2}-d^{2}\right)}.$
Figure 7. Locus of point $(x^{\prime},y^{\prime})$.
The coordinates of point $L_{1}(x_{3},y_{3})$ are $x_{3}=x_{0}+x_{1}-a$, $y_{3}=y_{0}-b$. Similarly, the coordinates of point $M_{1}(x_{4},y_{4})$ are $x_{4}=x_{0}+x_{2}-c$, $y_{3}=y_{0}-d$. Then the equations of lines $L_{1}K_{1}$ and $M_{1}K_{2}$ are $\frac{y-y_{3}}{x-x_{3}}=\frac{x_{1}-a}{b}$ and $\frac{y-y_{4}}{x-x_{4}}=\frac{x_{2}-c}{d}$, respectively. By solving the last two equations for $y$ we calculate ordinate $y^{\prime}$ of intersection point of lines $L_{1}K_{1}$ and $M_{1}K_{2}$ as
$y^{\prime}=\phi(t):=\frac{(bd-t^{2})\left(\sqrt{1-\frac{b^{2}}{t^{2}}}+\sqrt{1-\frac{d^{2}}{t^{2}}}\right)+\frac{bd(c-a)}{t}}{b\sqrt{1-\frac{d^{2}}{t^{2}}}+d\sqrt{1-\frac{b^{2}}{t^{2}}}},$
where the denominator is always positive and the numerator is a decreasing function for $t\geq b$. The abscissa $x^{\prime}$ of this intersection point can also be determined but we will not need it:
$x^{\prime}=-\frac{\left(b^{2}d-dt^{2}+b\sqrt{t^{2}-d^{2}}\,\sqrt{t^{2}-b^{2}}\right)}{\left(b^{2}-t^{2}\right)t^{2}\left(b^{2}-d^{2}\right)}\times$
×$\times\left(\left(ab\sqrt{t^{2}-b^{2}}+\left(b^{2}-t^{2}\right)\left(b-d\right)\right)\sqrt{t^{2}-d^{2}}\\ +t^{2}\left(b-d\right)\sqrt{t^{2}-b^{2}}-cd\left(b^{2}-t^{2}\right)\right).$
The formula for $x^{\prime}$ and $y^{\prime}$ determine a parametric curve which is locus of intersection point of lines $L_{1}K_{1}$ and $M_{1}K_{2}$. This locus can intersect line $DE$ at maximum 1 point for $t\geq b$. Indeed, note that $\phi(b)=d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}$ and $\phi(+\infty)=-\infty$. Therefore, if $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}\geq 0$, then equation $\phi(t)=0$ has a unique solution in interval $t\geq b$. Otherwise, if $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}<0$, then equation $\phi(t)=0$ has no solutions in interval $t\geq b$. In the statement of the problem we are given acute angled triangle $CDE$ satisfying all the conditions for extrema. Therefore, $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$ and the only solution of equation $\phi(t)=0$ in interval $t>b$ is $t_{0}=|AD|$. Note that $t_{0}^{2}=d^{2}+\lambda^{2}$, where $\lambda$ is one of the solutions of fourth degree equation
$\left(b^{2}-d^{2}\right)x^{4}+\left(2bda-2bc\right)x^{3}+\left(-2b^{3}d+2b^{2}d^{2}+2b\,d^{3}-2d^{4}\right)x^{2}$
$+\left(-2a\,b^{2}d^{2}+2ab\,d^{3}+2b^{2}cd-2bc\,d^{2}\right)x$
$+a^{2}b^{2}d^{2}+b^{4}d^{2}-2b^{3}d^{3}+2b\,d^{5}-d^{6}-2a\,b^{2}cd+b^{2}c^{2}=0,$
and therefore can be solved in radicals using Ferrari’s method.
It remains to show that if $b>d$ and $k_{2}>0$, then
$\lim_{t\rightarrow 0^{+}}f(a+k_{1}t,b+k_{2}t)>|CD|+|CE|.$
Since the limit is minimal when $k_{1}=0$, it is sufficient to show that
$\lim_{y\rightarrow b^{+}}f(a,y)>|CD|+|CE|.$
Note that since $2d^{3}+y^{3}\geq 3d^{2}y$,
$f_{yy}(a,y)=\frac{\left(a-c\right)^{2}\left(2d^{3}+{y^{3}}-{3d^{2}y}+2d\left(a-c\right)^{2}\right)}{\sqrt{\left(\left(y-d\right)^{2}+\left(a-c\right)^{2}\right)^{3}}\,\left(y-d\right)^{3}}>0,$
and therefore $f(a,y)$ is convex in interval $y\geq b$. Also $\lim_{y\rightarrow+\infty}f_{y}(a,y)=2>0$ and
$\lim_{y\rightarrow b^{+}}f_{y}(a,y)=1+\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}-\frac{b\left(a-c\right)^{2}}{\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}\,\left(b-d\right)^{3}}<0.$
Indeed,
$\lim_{y\rightarrow b^{+}}f_{y}(a,y)=\left(1+\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}\right)\left(1-\frac{d}{b-d}\left(\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}-1\right)\right),$
where the first factor is always positive. The second factor can be simplified as
$\frac{b-d\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}}{b-d}=\frac{b(b-d)-d\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}}{\left(b-d\right)^{2}}$
$=\frac{(b^{2}-d^{2})(b-d)^{2}-d^{2}(a-c)^{2}}{(b-d)^{2}\left(b(b-d)+d\sqrt{\left(a-c\right)^{2}+\left(b-d\right)^{2}}\right)},$
where the denominator is always positive. The numerator can be factorised as
$\left((b-d)\sqrt{b^{2}-d^{2}}-d(a-c)\right)\left((b-d)\sqrt{b^{2}-d^{2}}+d(a-c)\right),$
which is negative because as was mentioned before $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$. Therefore $f(a,y)$ is decreasing and then increasing for $y>b$ and consequently, $f(a,y)$ can not have infimum at $(a,b)$ and the only minimum of $f(x,y)$ is point $C$. ∎
As an immediate application of Theorem 1 we obtain the following particular result.
Corollary 2.2.
Let $CDE$ be a triangle with acute angles at vertices $D$ and $E$, such that $|CE|>|CD|$. Let $CK$ be the angle bisector of triangle $CDE$. Drop perpendiculars $KL$ and $KM$ to sides $CD$ and $CE$, respectively. Take points $A$ and $B$ on sides $CD$ and $CE$, respectively, such that $|AC|=|LD|$, $|BC|=|ME|$. Drop perpendicular $AA_{0}$, $BB_{0}$, and $CC_{0}$ to side $DE$. Extend line $AB$ to intersect line $DE$ at point $P$. Then $|AA_{0}|+|AP|>|CD|+|CE|$ (see Figure 9).
Note that we also solved the following more general problem.
Problem 1.
Given $A(a,b),\ B(c,d),$ with $a<c$ and $b\geq d>0$ find the minimum of $|C_{1}D_{1}|+|C_{1}E_{1}|$, where $C_{1}(x,y)$ is a point such that $y>b$ and rays $C_{1}A$ and $C_{1}B$ intersect line $y=0$ at points $D_{1}$ and $E_{1}$, respectively.
Answer. If $b=d$ or if $b>d$ and $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$, then the minimum exists and it is equal to $|CD|+|CE|$, from the previous result. If $b>d$ and $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}\leq 0$, then the minimum does not exist. In this case there is infimum of $|C_{1}D_{1}|+|C_{1}E_{1}|$ at point $(a,b)$ which is equal to $b\left(1+\sqrt{1+\frac{\left(a-c\right)^{2}}{\left(b-d\right)^{2}}}\right)$. See Figure 8 for the graph of $f(x,y)$ in these two cases.
Figure 8. Graph of $f$ in cases $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}\leq 0$ (left, $a=0.1$, $b=0.3$, $c=0.3$, $d=0.1$) and $d-b+\frac{d(c-a)}{\sqrt{b^{2}-d^{2}}}>0$ (right, $a=0.1$, $b=0.3$, $c=0.3$, $d=0.2$). Black point $\left(a,b,\lim_{y\rightarrow b^{+}}{f(a,y)}\right)$.
Remark 1. Note that in general $f(x,y)$ is not a convex function. So, we can not use the well know result in analysis which says that a local minimum of a convex function is also the absolute minimum (see e.g. [10], p. 543). Actually, each of functions
$f_{1}\left(x,y\right)=y\cdot\sqrt{1+\frac{\left(x-a\right)^{2}}{\left(y-b\right)^{2}}},\ f_{2}\left(x,y\right)=y\cdot\sqrt{1+\frac{\left(x-c\right)^{2}}{\left(y-d\right)^{2}}},$
have negative Hessian. Indeed,
$(f_{1})_{xx}(f_{1})_{yy}-(f_{1})_{xy}(f_{1})_{yx}=-\frac{b^{2}\left(x-a\right)^{2}}{\left(y-b\right)^{4}\left(\left(x-a\right)^{2}+(y-b)^{2}\right)}<0,$
and similarly for $f_{2}$. As some numerical examples show there can be points $(x,y)$ for which Hessian of $f=f_{1}+f_{2}$ is also negative.
Remark 2. We proved that triangle $CDE$ has optimal $|CD|+|CE|$ within all other triangles $C_{1}D_{1}E_{1}$, which share common pair of points $A$ and $B$ on their sides. Triangle $CDE$ has also other interesting properties worth mentioning. First drop perpendiculars $AA_{0}$, $BB_{0}$, and $CC_{0}$ to side $DE$. Extend lines $AB$ and $ML$ to intersect line $DE$ at points $P$ and $Q$, respectively (see Figure 9).
Figure 9. Properties of $\triangle CDE$.
  • $|PE|=|QD|=\frac{|DE|^{2}}{|B_{0}C_{0}|-|A_{0}C_{0}|}$. In particular, this implies that if points $D$, $E$, and $C_{0}$ are fixed, and point $C$ moves vertically along $CC_{0}$, then lines $AB$ and $LM$ pass though fixed points $P$ and $Q$, respectively.
  • $|A_{0}B_{0}|=\frac{[CDE]}{R}+|DE|$, where $[CDE]$ and $R$ are the area and the circumradius, respectively, of $\triangle CDE$.
  • $\frac{|AA_{0}|}{|BB_{0}|}=\frac{|CE|}{|CD|}$.
  • $|CA|+|CB|=\frac{|DE|^{2}}{|CE|+|CD|}$.
  • $|PQ|=\frac{|DE|^{3}}{|CE|^{2}-|CD|^{2}}$.
  • $\frac{|CA|}{|CB|}=\frac{|DC_{0}|}{|EC_{0}|}$.
Remark 3. I leave as an open question to prove that if $CK$ in Theorem 2.1 is a symedian, then $|C_{1}D_{1}|^{2}+|C_{1}E_{1}|^{2}>|CD|^{2}+|CE|^{2}$. As a problem for further exploration it would also be interesting to study these questions for more than two points instead of just two points $A$ and $B$. One can also replace line $DE$ with a plane and ask the same question for more than two points in space.
3. Conclusion
In the paper we solved a geometric optimization problem inspired by Philo’s line. The obtained solution reveals interesting properties of the angle bisector and its projections on the sides of this triangle. The provided proof uses multivariable optimization and geometrical methods. An open problem about symedians is proposed.
Acknowledgments
4. Declarations
Ethical Approval. Not applicable.
Competing interests. None.
Authors’ contributions. Not applicable.
Funding. This work was completed with the support of ADA University Faculty Research and Development Fund.
Availability of data and materials. Not applicable